The power required to lift a certain body through a distance of 30m in 6 seconds is 200 watts. Calculate workdone and mass of the body.
1
Expert's answer
2019-12-13T09:53:03-0500
"\rm work\: done=power\times time""W=Pt=200\:\rm W\times 6 \:s=1200\: J"

Also we have


"W=mgh"

So, the mass of the body

"m=\frac{W}{gh}=\frac{1200\:\rm J}{10\rm N/kg\times 30\: m}=4\: \rm kg"